# Advancing Semantic Future Prediction through Multimodal Visual Sequence Transformers

### v1

@@ Abstract
Semantic future prediction is important for autonomous systems navigating dynamic environments. This paper introduces FUTURIST, a method for multimodal future semantic prediction that uses a unified and efficient visual sequence transformer architecture. Our approach incorporates a multimodal masked visual modeling objective and a novel masking mechanism designed for multimodal training. This allows the model to effectively integrate visible information from various modalities, improving prediction accuracy. Additionally, we propose a VAE-free hierarchical tokenization process, which reduces computational complexity, streamlines the training pipeline, and enables end-to-end training with high-resolution, multimodal inputs. We validate FUTURIST on the Cityscapes dataset, demonstrating state-of-the-art performance in future semantic segmentation for both short- and mid-term forecasting. We provide the implementation code at https://github.com/Sta8is/FUTURIST.
@@ 2 Related Work
Future Video Forecasting  Future video forecasting and generation have garnered significant attention [48] in the computer vision and machine learning communities due to their potential applications in various domains, including autonomous driving and robotics. Early approaches to this task often relied on recurrent neural networks (RNNs) such as Convolutional Long Short Term Memory (Conv-LSTM) architectures [46, 20, 67, 70, 65, 37, 8] to capture temporal dependencies in video sequences. However, these methods frequently struggled with long-term predictions and maintaining visual consistency and realism. Other works have leveraged generative adversarial networks (GANs) and variational autoencoders (VAEs) [64, 36, 8, 71, 2] to produce more realistic and diverse future frames. More recently, techniques leveraging diffusion models [26, 27, 19], demonstrate impressive visual results capturing both spatial and temporal relationships. Finally, inspired from their success in natural language models, some works extend the ideas of auto-regressive and masked modeling objectives to images [9, 10] and videos [66, 23, 75, 76] and employ pure transformer models in conjunction with a tokenization module [52, 62, 18]. In our work, we propose a method that directly predicts semantic modalities for future frames without the need for complex tokenization modules, offering a simpler and more efficient approach to future forecasting.
Multimodal Learning with Transformers  To effectively process multiple dense visual modalities and enable robust cross-modal interactions, several multimodal transformer approaches have been proposed. Some approaches employ flexible architectures that handle arbitrary input and output modalities. Notably, Perceiver IO [31] and ImageBind [21] integrate embeddings from multiple modalities into a shared representation space, enabling unified processing of diverse data types within a single framework. Other works extend masked autoencoders to multimodal and multitask settings, such as MultiMAE [3], which introduces a masked transformer for joint pretraining across modalities. Building upon this, models for massively multimodal masked modeling and any-to-any vision tasks have been introduced [45, 4].
@@
Semantic Future Prediction  Future segmentation and depth prediction aim to use past frames to predict the segmentation masks and depth maps of future frames. In semantic segmentation, Luc et al. [42] introduced the first method for predicting future semantics by exploring different input configurations. They experimented with using previous RGB frames, past segmentation maps, or a combination of both. Their results showed that a two-phase approach—predicting future frame pixel values and then using these predictions to generate segmentation masks—was less effective. We further validate this finding by applying recent diffusion-based, state-of-the-art driving world models for future RGB frame generation.
To better capture the temporal nature of videos, several approaches have been proposed. Focusing on the architecture, [46, 58] integrated ConvLSTM modules, while [11] introduced an attention mechanism. [12] adopted a teacher-student framework, where the student performs the forecasting task, and the teacher provides semantic guidance through knowledge distillation. Another promising direction involves flow-based forecasting [32, 60, 55], where the prediction model uses optical flow information. Jin et al. [32] developed a method that predicts both optical flow and future semantic segmentations by allowing the segmentation model to leverage motion features from a flow-predicting network. Terwilliger et al. [60] further improved this by using a convolutional LSTM to predict optical flow and then warp the last segmentation mask. In a similar approach, Ciamarra et al. [13] applied optical flow warping to instance segmentation forecasting.
In recent work, [55] proposed the F2MF model, which extends the F2F (Feature-to-Feature) approach [43], where features are regressed from observed frames. F2MF improves upon this by warping deep features to generate future semantic segmentation. For future depth prediction, Liu et al. [40, 47] tackled the problem of predicting depth maps for unobserved frames. Depth prediction was also explored in [28, 49, 14] as part of multimodal prediction, where depth was combined with other modalities like optical flow, semantic segmentation, and scene understanding to enhance the accuracy and robustness of future scene interpretations.
@@ 3 Multimodal Visual Sequence Transformer
We propose FUTURIST, a flexible framework for semantic future prediction that processes video sequences in a spatially dense, multimodal way. Our framework uses a video sequence of $N_{c}$ frames from timesteps $\{t-N_{c}+1,\ldots,t\}$ as context for making predictions. These input frames may represent any semantic modality, such as semantic segmentation or depth maps, or multiple modalities simultaneously. The objective is to predict future frames for the input modalities in a frame-wise autoregressive manner, with a focus on the short-term (${t+1}$) and mid-term (${t+3}$) future, consistent with previous studies [42, 43, 12, 39].
@@
Pseudo-Labeled Multimodal Training Dataset
@@ 3.1 VAE-free Multimodal Embedder
Integrated Per-Modality Tokenization Process
@@
Efficient Multimodality Fusion
@@ 3.2 Multimodal Masking Strategies
[…]
Inference Time
@@ 3.5 End-to-End Training for Masked Multimodal Visual Modeling
Per-Modality Masked Modeling Loss
@@
where $w_{\mathcal{I}_{i}}$ is the weight coefficient for the loss of the $\mathcal{I}_{i}$ modality. Unlike most transformer-based generative models that rely on pre-trained tokenizers, our framework trains all modules—the embedders (Sec. 3.1), the masked transformer backbone (Sec. 3.3), and the decoder heads (Sec. 3.4)—together using the objective defined in Eq. 2. This end-to-end approach ensures that the entire model is optimized jointly, making our multimodal visual modeling framework both simple and effective.
@@ 4.1 Experimental Protocol
Evaluation Metrics.  To assess our method’s performance in predicting future semantic segmentation and depth maps, we use the following evaluation metrics.
@@
Evaluation Scenarios  Following prior work [42, 46], we consider two evaluation scenarios: short-term prediction (3 frames ahead,  0.18s) and mid-term prediction (9 frames ahead,  0.54s). In both cases, the 20th frame is the forecasting target. We temporally subsample sequences by a factor of 3 before processing them with our model. This means that, for short-term prediction, a model with a context length of four frames (i.e., $N_{c}=4$ and $N_{p}=1$) uses frames 8, 11, 14, and 17 as context to predict frame 20. For mid-term prediction, the model uses frames 2, 5, 8, and 11 as context and predicts frame 20 auto-regressively (first frame 14, then 17, and finally frame 20).
@@ 4.2 Quantitative Results
Baselines  We compare our method against three baselines. The first is the Oracle baseline, which applies the image segmentation model (Segmenter [57]) or the monocular depth prediction model (DepthAnythingV2 [74]) directly on the target future frame. Since these models are the same ones used to compute segmentation and depth maps for the context frames, the Oracle baselines serve as an upper bound for our method’s performance.
@@
The third baseline uses the state-of-the-art driving world model, VISTA [19], which generates future RGB frames conditioned on three context frames through a video diffusion model. VISTA is a large-scale model with 2.5 billion parameters, trained on 1,740 hours of driving videos from YouTube. In this baseline, we use VISTA (with action-free conditioning) to generate future RGB frames, then apply the Segmenter and DepthAnythingV2 models to these frames to produce future-frame segmentation and depth maps. This comparison allows us to assess our method, which directly predicts future segmentation and depth maps, against the state-of-the-art in generative-based future prediction.
@@
Comparison with Prior Work on Segmentation Forecasting  In Table 1, we compare our model with previous semantic future prediction methods on the Cityscapes dataset. We report results for both the default multimodal version (using segmentation and depth modalities) and a segmentation-only variant. Both versions outperform prior methods in short- and mid-term predictions, with the multimodal version achieving the best performance and setting a new state-of-the-art. Our model shows significant improvements over the previous state-of-the-art method, PFA [39], especially for movable object (MO) classes, which are typically more difficult to predict.
@@
Comparison with VISTA  We also provide results for the VISTA-based baseline in Table 1. VISTA performs notably worse, even lower than the Copy-Last baseline. As shown in Figure 5, this is due to VISTA’s inability to maintain the semantic structure of the scene in future frames. Despite generating visually realistic frames, VISTA’s future predictions often suffer from issues such as object disappearance and artifacts that lead to inaccurate results from the oracle segmentation and depth models. These limitations highlight the advantage of our approach, which models future predictions directly in semantic space. Complete quantitative results, including depth metrics, are provided in Appendix Tab. 6.
Multi-modal vs Single-modality models  In Table 2, we compare our model’s default multimodal version (segmentation plus depth) to single-modality versions using only segmentation or only depth. Results for both segmentation and depth forecasting show that the multimodal model outperforms the single-modality versions, demonstrating that it captures synergistic effects between the two modalities, enabling more accurate future predictions.
@@
Table 5: Sequence Length Impact. Models trained for 800 epochs with sequence lengths $N$ = 4, 5, and 7, with $N_{p}=1$.
Scaling Training Epochs  Building on research into scaling laws for transformer-based architectures [33, 77], we investigate how our model’s performance scales with the number of training epochs. As shown in Figure 4, extending the training duration results in significant improvements in both segmentation and depth prediction. These results suggest that our model can continue to enhance its predictive capabilities with more training, even when trained on limited datasets like Cityscapes.
@@
Masking Strategies  In Table 4, we compare various multimodal training masking strategies discussed in Section 3.2. The Partially-shared + Exclusive strategy achieves the best results across all metrics. This strategy restricts future-frame masking during training such that each future-frame token is either fully masked (both modalities are hidden) or partially masked (only one modality is visible). Importantly, no future-frame token has both modalities visible at the same time. We believe this approach helps the model better learn the synergistic effects between the two modalities. In contrast, the Fully-Masked strategy yields the worst results, despite matching the masking approach used during inference. We hypothesize this is because the strategy is too aggressive and less effective in promoting the model’s learning of interactions between the modalities.
Sequence Length Impact  In Table 5 we compare models with sequence lengths $N$ of 4, 5, and 7 frames (with $N_{p}=1$). While all models perform similarly, the 7-frame model achieves slightly better results for mid-term segmentation and all depth prediction metrics. However, we use the 5-frame context by default in this paper due to the higher computational cost of the 7-frame model.
Additional Experiments  In the Appendix A, we present additional experiments validating our design choices. First, in Appendix Tab. 7 we demonstrate that our VAE-free tokenization approach significantly outperforms alternative tokenization methods using pretrained VQ-VAE models, particularly for semantic modalities like segmentation where VQ-VAE reconstruction severely degrades performance. Second, in Appendix Tab. 8 we examine our multi-modal fusion strategy against a separable tokens approach, showing that under equal compute budgets, our method achieves superior performance on most metrics.
@@ 5 Conclusion
We evaluated FUTURIST on the Cityscapes dataset for semantic segmentation and depth prediction. Our results show accurate short-term and mid-term predictions, outperforming single-modality models and setting a new state-of-the-art in future semantic segmentation. These findings demonstrate FUTURIST’s effectiveness in extending masked transformers for multimodal future prediction. We also show that FUTURIST scales well with longer training times, even on Cityscapes. Future work includes scaling to larger datasets and adding new modalities.
Despite its strengths, FUTURIST lacks action conditioning, limiting controllability in autonomous driving, and may miss subtle visual details, such as slight road surface variations, due to its focus on semantic prediction. For a comprehensive discussion of limitations and future directions, see also Appendix C.
@@ Appendix A Additional Results
Copy Last 55.6 52.7 40.5 32.3 90.5 10.780 82.2 18.345
VISTA [19] 45.7 43.8 40.4 35.6 84.9 14.325 80.2 18.335
@@
A.1 Comparison with VISTA
In Table 6 we compare our FUTURIST approach against the VISTA [19] baseline on both the segmentation and depth forecasting tasks. FUTURIST achieves clearly better results.
A.2 Tokenization: Comparing Our VAE-Free Approach to VQ-VAE
@@
The main reason for this gap is that the reconstruction process in VQ-VAE significantly degrades the oracle performance, particularly for the segmentation modality. In fact, segmentation results after VQ-VAE reconstruction are worse than our predicted segmentation results for short-term predictions. While retraining VQ-VAE models specifically for the considered modalities might mitigate this issue, such retraining requires substantial effort and falls outside the scope of this work.
The key takeaway is that while VQ-VAE tokenizers are essential for generative models focused on image or video generation, they are likely unnecessary for semantic modalities like those considered here. Our VAE-free approach not only simplifies the training pipeline but also achieves superior performance
A.3 Impact of Our Multi-Modal Fusion Strategy

### v2

@@ Abstract
Semantic future prediction is important for autonomous systems navigating dynamic environments. This paper introduces FUTURIST, a method for multimodal future semantic prediction that uses a unified and efficient visual sequence transformer architecture. Our approach incorporates a multimodal masked visual modeling objective and a novel masking mechanism designed for multimodal training. This allows the model to effectively integrate visible information from various modalities, improving prediction accuracy. Additionally, we propose a VAE-free hierarchical tokenization process, which reduces computational complexity, streamlines the training pipeline, and enables end-to-end training with high-resolution, multimodal inputs. We validate FUTURIST on the Cityscapes dataset, demonstrating state-of-the-art performance in future semantic segmentation for both short- and mid-term forecasting. Project page and code at https://futurist-cvpr2025.github.io/.
@@ 2 Related Work
Future Video Forecasting.  Future video forecasting and generation have garnered significant attention [50] in the computer vision and machine learning communities due to their potential applications in various domains, including autonomous driving and robotics. Early approaches to this task often relied on recurrent neural networks (RNNs) such as Convolutional Long Short Term Memory (Conv-LSTM) architectures [48, 20, 73, 76, 71, 39, 8] to capture temporal dependencies in video sequences. However, these methods frequently struggled with long-term predictions and maintaining visual consistency and realism. Other works have leveraged generative adversarial networks (GANs) and variational autoencoders (VAEs) [70, 38, 8, 77, 2] to produce more realistic and diverse future frames. More recently, techniques leveraging diffusion models [28, 29, 19], demonstrate impressive visual results capturing both spatial and temporal relationships. Finally, inspired from their success in natural language models, some works extend the ideas of auto-regressive and masked modeling objectives to images [9, 10] and videos [72, 25, 81, 82] and employ pure transformer models in conjunction with a tokenization module [57, 68, 18]. In our work, we propose a method that directly predicts semantic modalities for future frames without the need for complex tokenization modules, offering a simpler and more efficient approach to future forecasting.
Multimodal Learning with Transformers.  To effectively process multiple dense visual modalities and enable robust cross-modal interactions, several multimodal transformer approaches have been proposed. Some approaches employ flexible architectures that handle arbitrary input and output modalities. Notably, Perceiver IO [33] and ImageBind [21] integrate embeddings from multiple modalities into a shared representation space, enabling unified processing of diverse data types within a single framework. Other works extend masked autoencoders to multimodal and multitask settings, such as MultiMAE [3], which introduces a masked transformer for joint pretraining across modalities. Building upon this, models for massively multimodal masked modeling and any-to-any vision tasks have been introduced [47, 4].
@@
Semantic Future Prediction.  Future segmentation and depth prediction aim to use past frames to predict the segmentation masks and depth maps of future frames. In semantic segmentation, Luc et al. [44] introduced the first method for predicting future semantics by exploring different input configurations. They experimented with using previous RGB frames, past segmentation maps, or a combination of both. Their results showed that a two-phase approach—predicting future frame pixel values and then using these predictions to generate segmentation masks—was less effective. We further validate this finding by applying recent diffusion-based, state-of-the-art driving world models for future RGB frame generation.
To better capture the temporal nature of videos, several approaches have been proposed. Focusing on the architecture, [48, 64] integrated ConvLSTM modules, while [11] introduced an attention mechanism. [12] adopted a teacher-student framework, where the student performs the forecasting task, and the teacher provides semantic guidance through knowledge distillation. Another promising direction involves flow-based forecasting [34, 66, 60, 61, 13], where the prediction model uses optical flow information. Jin et al. [34] developed a method that predicts both optical flow and future semantic segmentations by allowing the segmentation model to leverage motion features from a flow-predicting network. Terwilliger et al. [66] further improved this by using a convolutional LSTM to predict optical flow and then warp the last segmentation mask. In a similar approach, Ciamarra et al. [13] applied optical flow warping to instance segmentation forecasting.
In recent work, [60, 61] proposed the F2MF model, which extends the F2F (Feature-to-Feature) approach [45], where features are regressed from observed frames, by warping deep features to generate future semantic segmentation. For instance and panoptic segmentation forecasting, several methods [23, 54, 24] have been proposed that extend beyond semantic boundaries to predict individual object instances. For future depth prediction, Liu et al. [42, 49] tackled the problem of predicting depth maps for unobserved frames. Depth prediction was also explored in [30, 53, 14] as part of multimodal prediction, where depth was combined with other modalities like optical flow, semantic segmentation, and scene understanding to enhance the accuracy and robustness of future scene interpretations.
@@ 3 Multimodal Visual Sequence Transformer
We propose FUTURIST, a flexible framework for semantic future prediction that processes video sequences in a spatially dense, multimodal way. Our framework uses a video sequence of $N_{c}$ frames from timesteps $\{t-N_{c}+1,\ldots,t\}$ as context for making predictions. These input frames may represent any semantic modality, such as semantic segmentation or depth maps, or multiple modalities simultaneously. The objective is to predict future frames for the input modalities in a frame-wise autoregressive manner, with a focus on the short-term (t+1{t+1}) and mid-term (t+3{t+3}) future, consistent with previous studies [44, 45, 12, 41].
@@
Pseudo-Labeled Multimodal Training Dataset.
@@ 3.1 VAE-free Multimodal Embedder
Integrated Per-Modality Tokenization Process.
@@
Efficient Multimodality Fusion.
@@ 3.2 Multimodal Masking Strategies
For each modality $\mathcal{I}$, the future-frame tokens in $\mathbf{Z}_{\mathcal{I}}$ indicated as masked in $\mathbf{M}_{\mathcal{I}}$ are replaced with a modality-specific learnable mask embedding $\textsc{[mask]}_{\mathcal{I}}\in\mathbb{R}^{d_{\mathcal{I}}}$.
Inference Time.
@@ 3.5 End-to-End Training for Masked Multimodal Visual Modeling
Per-Modality Masked Modeling Loss.
@@
where $w_{\mathcal{I}_{i}}$ is the weight coefficient for the loss of the $\mathcal{I}_{i}$ modality. Unlike most transformer-based generative models that rely on pre-trained tokenizers, our framework trains all modules—the embedders (subsection 3.1), the masked transformer backbone (subsection 3.3), and the decoder heads (subsection 3.4)—together using the objective defined in Equation 2. This end-to-end approach ensures that the entire model is optimized jointly, making our multimodal visual modeling framework both simple and effective.
@@ 4.1 Experimental Protocol
Evaluation Metrics.  To assess our method’s performance in predicting future semantic segmentation and depth maps, we evaluate on full-resolution using the following evaluation metrics.
@@
Evaluation Scenarios.  Following prior work [44, 48], we consider two evaluation scenarios: short-term prediction (3 frames ahead,  0.18s) and mid-term prediction (9 frames ahead,  0.54s). In both cases, the 20th frame is the forecasting target. We temporally subsample sequences by a factor of 3 before processing them with our model. This means that, for short-term prediction, a model with a context length of four frames (i.e., $N_{c}=4$ and $N_{p}=1$) uses frames 8, 11, 14, and 17 as context to predict frame 20. For mid-term prediction, the model uses frames 2, 5, 8, and 11 as context and predicts frame 20 auto-regressively (first frame 14, then 17, and finally frame 20).
@@ 4.2 Quantitative Results
Baselines.  We compare our method against three baselines. The first is the Oracle baseline, which applies the image segmentation model (Segmenter [63]) or the monocular depth prediction model (DepthAnythingV2 [80]) directly on the target future frame. Since these models are the same ones used to compute segmentation and depth maps for the context frames, the Oracle baselines serve as an upper bound for our method’s performance.
@@
The third baseline uses the state-of-the-art driving world model, VISTA [19], which generates future RGB frames conditioned on three context frames through a video diffusion model. VISTA is a large-scale model with 2.5 billion parameters, trained on 1,740 hours of driving videos from YouTube. In this baseline, we finetune VISTA on Cityscapes and use it (with action-free conditioning) to generate future RGB frames, then apply the Segmenter and DepthAnythingV2 models to these frames to produce future-frame segmentation and depth maps. This comparison allows us to assess our method, which directly predicts future segmentation and depth maps, against the state-of-the-art in generative-based future prediction. More details regarding Vista finetuning in Appendix A.1.
@@
Comparison with Prior Work on Segmentation Forecasting.  In Table 1, we compare our model with previous semantic future prediction methods on the Cityscapes dataset. We report results for both the default multimodal version (using segmentation and depth modalities) and a segmentation-only variant. Both versions outperform prior methods in short- and mid-term predictions, with the multimodal version achieving the best performance and setting a new state-of-the-art. Our model shows significant improvements over the previous state-of-the-art method, PFA [41], especially for movable object (MO) classes, which are typically more difficult to predict.
@@
Comparison with VISTA.  We also provide results for the VISTA-based baseline in Table 1. VISTA performs notably worse, even lower than the Copy-Last baseline. As shown in Figure 5, this is due to VISTA’s inability to maintain the semantic structure of the scene in future frames. Despite generating visually realistic frames, VISTA’s future predictions often suffer from issues such as object disappearance and artifacts that lead to inaccurate results from the oracle segmentation and depth models. These limitations highlight the advantage of our approach, which models future predictions directly in semantic space. Complete quantitative results, including depth metrics, are provided in Appendix Tab.  6.
Multi-modal vs Single-modality models.  In Table 2, we compare our model’s default multimodal version (segmentation plus depth) to single-modality versions using only segmentation or only depth. Results for both segmentation and depth forecasting show that the multimodal model outperforms the single-modality versions, demonstrating that it captures synergistic effects between the two modalities, enabling more accurate future predictions.
@@
Table 5: Sequence Length Impact. Models trained for 800 epochs with sequence lengths NN = 4, 5, and 7, with $N_{p}=1$.
Scaling Training Epochs.  Building on research into scaling laws for transformer-based architectures [35, 83], we investigate how our model’s performance scales with the number of training epochs. As shown in Figure 4, extending the training duration results in significant improvements in both segmentation and depth prediction. These results suggest that our model can continue to enhance its predictive capabilities with more training, even when trained on limited datasets like Cityscapes.
@@
Masking Strategies.  In Table 4, we compare various multimodal training masking strategies discussed in subsection 3.2. The Partially-shared + Exclusive strategy achieves the best results across all metrics. This strategy restricts future-frame masking during training such that each future-frame token is either fully masked (both modalities are hidden) or partially masked (only one modality is visible). Importantly, no future-frame token has both modalities visible at the same time. We believe this approach helps the model better learn the synergistic effects between the two modalities. In contrast, the Fully-Masked strategy yields the worst results, despite matching the masking approach used during inference. We hypothesize this is because the strategy is too aggressive and less effective in promoting the model’s learning of interactions between the modalities.
Sequence Length Impact.  In Table 5 we compare models with sequence lengths $N$ of 4, 5, and 7 frames (with $N_{p}=1$). While all models perform similarly, the 7-frame model achieves slightly better results for mid-term segmentation and all depth prediction metrics. However, we use the 5-frame context by default in this paper due to the higher computational cost of the 7-frame model.
Additional Experiments.  In the Appendix A, we present additional experiments validating our design choices. First, in Appendix Tab. 7 we demonstrate that our VAE-free tokenization approach significantly outperforms alternative tokenization methods using pretrained and finetuned VQ-VAE models. Second, in Appendix Tab. 8 we examine our multi-modal fusion strategy against a separable tokens approach, showing that under equal compute budgets, our method achieves superior performance on most metrics.
@@ 5 Conclusion
We evaluated FUTURIST on the Cityscapes dataset for semantic segmentation and depth prediction. Our results show accurate short-term and mid-term predictions, outperforming single-modality models and setting a new state-of-the-art in future semantic segmentation. These findings demonstrate FUTURIST’s effectiveness in extending masked transformers for multimodal future prediction. We discuss the limitations of our method and directions for future work in Appendix C.
@@ Appendix A Additional Results
Copy Last 55.5 52.7 40.5 32.2 90.5 10.780 82.2 18.345
VISTA Fine-tuned 65.7 64.1 53.2 49.1 93.4    8.134 87.4 12.923
@@
A.1 Comparison with VISTA
In Table 6, we compare our FUTURIST approach against the VISTA [19] baseline on both the segmentation and depth forecasting tasks. We conducted evaluations using a version fine-tuned on Cityscapes for 10 epochs to address potential performance issues arising from domain shift. LORA Finetuning required 8 GPUs, with a batch size of 1 per GPU, utilizing approximately $80\times 8=640$GB of VRAM. VISTA finetuned still fell far behind FUTURIST.
A.2 Tokenization: Comparing Our VAE-Free Approach to VQ-VAE
The main reason for this gap is that the reconstruction process in VQ-VAE significantly degrades the oracle performance, particularly for the segmentation modality. In fact, segmentation results after VQ-VAE reconstruction are worse than our predicted segmentation results for short-term predictions. Moreover, we also fine-tuned LDM’s VQ-VAE [58] (since DALL-E does not offers training code and recipe) on segmentation and depth maps, improving both Oracle reconstruction (Oracle after LDM’s VQ-VAE-FT Reconstruction) and future segmentation prediction (VQ-VAE-FT from LDM), though depth prediction performance slightly declined. Still, our VAE-free method is more effective and efficient: fewer parameters (465M vs 558M total), faster training (8h vs 22h at 800 epochs), and lower inference time per sequence (52ms vs 274ms).
The key takeaway is that while VQ-VAE tokenizers are essential for generative models focused on image or video generation, they are likely unnecessary for semantic modalities like those considered here. Our VAE-free approach not only simplifies the training pipeline but also achieves superior performance.
A.3 Impact of Our Multi-Modal Fusion Strategy